OpenStudy (xximhisgirlxx):

Check my answers please? & tell me how to do it? 1. Find the slope. 2 –one half –2 one half

OpenStudy (xximhisgirlxx):

I think its D, could someone please explain to me how to find slopes. https://www.connexus.com/content/media/527934-512012-100107-AM-2077201970.jpg @amistre64

OpenStudy (anonymous):

Slope is \[\frac{ rise }{ run }\] so count up from the bottom point and then over to the top point. That will give you\[\frac{ 8 }{ -4 }\] Simplify and you get -2

OpenStudy (xximhisgirlxx):

will you please please help with more? I have my answers I just wanna make sure they are correct. @Novalynne

OpenStudy (anonymous):

Sure go ahead.

OpenStudy (xximhisgirlxx):

Use the graph below to answer the following question. 2. Find the slope of the line. Describe how one variable changes in relation to the other. https://www.connexus.com/content/media/527934-512012-101635-AM-2142815969.jpg 2; distance increases by 2 miles per hour 2; distance decreases by 2 miles per hour one half; distance increases 1 mile every 2 hours one half; distance decreases 1 mile every 2 hours

OpenStudy (xximhisgirlxx):

I think its A.

OpenStudy (xximhisgirlxx):

@Novalynne

OpenStudy (anonymous):

Correct.

OpenStudy (xximhisgirlxx):

Use the graph below to answer the following question. 3. Find the slope of the line. Describe how one variable changes in relation to the other. https://www.connexus.com/content/media/527934-512012-102129-AM-172483599.jpg two over three; the amount of water decreases by 2 gallons every 3 minutes. negative two thirds; the amount of water decreases by 2 gallons every 3 minutes. negative three halves; the amount of water decreases by 3 gallons every 2 minutes. –1 ; the amount of water decreases by 1 gallon per minute.

OpenStudy (xximhisgirlxx):

I think its B

OpenStudy (xximhisgirlxx):

@Novalynne

OpenStudy (anonymous):

Incorrect. The slope is \[\frac{ 1 }{ -1 }\] When you simplify you get -1. So it decreases by 1 gallon every minute.

OpenStudy (xximhisgirlxx):

4. The data in the table are linear. Use the table to find the slope. x|2|4|6|8 y|1|-2|-5|-8 https://www.connexus.com/content/media/103510-6302006-111010-AM-2027758843.gif

OpenStudy (xximhisgirlxx):

3over2 -3over2 -2over3 2over3

OpenStudy (xximhisgirlxx):

@Novalynne

OpenStudy (xximhisgirlxx):

I think its C

OpenStudy (xximhisgirlxx):

@Novalynne

OpenStudy (anonymous):

Actually it's B

OpenStudy (xximhisgirlxx):

@Novalynne

OpenStudy (xximhisgirlxx):

Its just this one and one more

OpenStudy (xximhisgirlxx):

7. Find the slope of a line that is parallel to the line containing the points (3, 4) and (2, 6). (1 point) m = 1 m = 2 m = –2 m = one half I think this one is C

OpenStudy (xximhisgirlxx):

are you there